OpenStudy (anonymous):

Help please --- Strep throat is caused by Streptococcus bacteria. The amount of bacteria multiplies at a rate of 10^13 per 10^2 hrs. What is the average growth of bacteria per hour?? Please help and explain the steps to answer. Thank you! (:

jigglypuff314 (jigglypuff314):

10^13 ------ = 10^2

OpenStudy (anonymous):

So thats how you find the average growth per hour??

jigglypuff314 (jigglypuff314):

growth per hour growth = ------- hour

OpenStudy (anonymous):

100000000000

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Thank you!! C:

jigglypuff314 (jigglypuff314):

glad I could help :)

OpenStudy (anonymous):

Usually bacteria grow exponentially. N(t) = N(0) exp(kt) ln[N(t)/N(0)] = k t ln[10^13/1] = 30 = k [100 hrs] k = 0.30/hr t in hours